find the dimensions of a rectangle whose width is 3 miles less than its length and whose area is 88 square miles.
Answer by amarjeeth123(569)

Let the length of the rectangle be x.
Width of the rectangle is (x-3).
Area of the rectangle=x(x-3).
x(x-3)=88
x^2-3x-88=0
x^2-11x+8x-88=0
x(x-11)+8(x-11)=0
(x-11)(x+8)=0
x=11.
The length of the rectangle is 11 miles and the width of the rectangle is 8 miles.